A cylindrical has a radius of 4.3 cm and a height of 8.1 cm. How much will a conical glass that has a radius of 8.6 cm and a height of 5.7 cm hold? Is this more or less than the cylindrical glass?
Answer by Alan3354(21581)

A cylindrical has a radius of 4.3 cm and a height of 8.1 cm. How much will a conical glass that has a radius of 8.6 cm and a height of 5.7 cm hold? Is this more or less than the cylindrical glass?
----------------
Cylinder volume:
V = pi*r^2*h
V = 4.3^2*8.1pi
V = 149.769pi cubic cm
------------------
Cone volume:
V = pi*r^2*h/3
V = 8.6^2*5.7pi/3
V = 140.524pi cubic cm
----------------
The cylinder is larger.